# Average-Case Deterministic Query Complexity of Boolean Functions with Fixed Weight

### v2

@@ 4.1 Upper bound
As a warm-up, we show a simple upper bound $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+O(1)$ for any boolean function $f$. Say a query algorithm is reasonable if it stops querying as long as the subfunction is constant. For convenience, we say input $x$ is a black point if $f(x)=1$.
@@
$\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+2$ for any non-constant boolean function $f$.
@@
Let $m=\operatorname{wt}(f)$. We prove Lemma 4.1 by induction on $m$. When $m=1$, we have $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})\leq 2$ by Proposition B.1. Assume the variable $x_{i}$ is queried first. Let $m_{0}=\operatorname{wt}(f|_{x_{i}=0})$ and $m_{1}=\operatorname{wt}(f|_{x_{i}=1})$. If $m_{b}=0$, where $b\in\{0,1\}$, any reasonable algorithm will stop on a constant subfunction $f|_{x_{i}=b}$. So the $\mathrm{D}_{\mathrm{ave}}(f)\leq 1+\frac{1}{2}(\log m+2)\leq\log m+2$ by the induction hypothesis. Otherwise, by the induction hypothesis and the AM-GM inequality, we have
@@
To prove an upper bound on $\mathrm{D}_{\mathrm{ave}}(f)$, it suffices to design a query algorithm computing $f$. The query algorithm we design for proving Lemma 4.1 and Theorem 1.1 are based on the same approach: query an arbitrary variable until the subfunction meets certain border conditions. The border conditions in the proofs:
  • Lemma 4.1: $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ if $\operatorname{wt}(f)=1$ (Proposition B.1).
  • Theorem 1.1: $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ if $\operatorname{wt}(f)\leq 4\log n$ (Corollary 4.5).
To prove the second one, let us fix the $\operatorname{wt}(f)$ and put the black points in a $\operatorname{wt}(f)\times n$ table with each row representing a black point. Notice the fixed number of distinct columns up to $2^{\operatorname{wt}(f)}$. When $n\gg 2^{\operatorname{wt}(f)}$, numerous duplicate columns are identified by the pigeonhole principle. Therefore, numerous black points share the same bit on multiple coordinates, leading to the construction of a decision tree with $O(1)$ average depth.
Suppose $f\in\mathcal{B}_{n,m}$ has $m$ black points $x^{(1)},\ldots,x^{(m)}\in\{0,1\}^{n}$. We call $c_{i}=(x_{i}^{(1)},\ldots,x_{i}^{(m)})$ the black column of the coordinate $i$. Say two coordinates $i,j$ are positively correlated if $c_{i}=c_{j}$ or negatively correlated if $c_{i}=\neg c_{j}$. An equivalent coordinate set (ECS) with respect to $f$ is a set of correlated coordinates. Say a coordinate set $S\subseteq\{1,\ldots,n\}$ is pure if $c_{i}$ is an all-zero or all-one vector for each $i\in S$; otherwise, say $S$ is mixed. For example, in Table 1, the set $\{5,9,{11}\}$ is pure, since $c_{5}=c_{9}=(0,0)$ and $c_{11}=(1,1)$, and the set $\{1,2,3\}$ is mixed, since $c_{1}=c_{3}=(0,1)$ and $c_{2}=(1,0)$.
@@
Fix a boolean function $f$. Suppose $i,j$ are positively correlated (respectively, negatively correlated). For every input $x\in\{0,1\}^{n}$, if $f(x)=1$, then $x_{i}=x_{j}$ (respectively, $x_{i}\neq x_{j}$).
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function, and $m=\operatorname{wt}(f)$. If $n>k\cdot 2^{m-1}$, there exists some maximal ECS of size $k+1$ at least.
Proposition 4.3 follows from the pigeonhole principle, since there are $2^{m-1}$ distinct maximal ECSs. Combining these facts, we can construct a “shallow” decision tree computing $f$ when $n$ is exponentially large.
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function. If $n>2^{\operatorname{wt}(f)}$, then $\mathrm{D}_{\mathrm{ave}}(f)\leq 5$.
@@
Let $m=\operatorname{wt}(f)$. We prove $\mathrm{D}_{\mathrm{ave}}(f)\leq 5$ by induction on $n$. By Proposition 4.3, there exists a maximal ECS $I=\{i_{1},\ldots,i_{k}\}$ for some $k\geq 3$. Assume $i_{1},\ldots,i_{k}$ are positively correlated without loss of generality. For any black point $x$, by Proposition 4.2, we have $x_{i_{1}}=\cdots=x_{i_{k}}$.
If $|I|=n$, any $f$’s black point $x$ satisfies $x_{1}=\cdots=x_{n}$. So there are at most two black points, all-zero or all-one. By Lemma 4.1, $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+2\leq 3$. Thus, we can assume $|I|<n$, and thus there exists a coordinate $j\not\in I$.
Case 1: The ECS $I$ is mixed. Without loss of generality, assume $c_{i_{1}}\in\{0,1\}^{m}$ is not all-zero or all-one. Our query algorithm $T_{1}$ runs as follows (see Figure 1(a)):
@@
    if $x_{i_{1}}=x_{i_{2}}$, query $x_{j}$, and apply an optimal query algorithm to the subfunction.
The decision tree $T_{1}$ computes $f$, since input $x$ cannot be a black point if $x_{i_{1}}\neq x_{i_{2}}$. Note that the probability of querying exactly 2 variables is $\frac{1}{2}$, and the probability of querying at least 3 variables is $\frac{1}{2}$.
(a) Case 1: The ECS $I$ is mixed and $|I|<n$
(b) Case 2: The ECS $I$ is pure and $|I|<n$
Figure 1: The decision trees with $O(1)$ average depth.
For convenience, denote by $\rho_{u,v}$ the restriction that fixes $x_{i_{1}},x_{i_{2}}$ to $u\in\{0,1\}$, fixes $x_{r}$ to $v\in\{0,1\}$, and leaves other variables undetermined. Since $c_{i_{1}}$ is neither all-zero nor all-one, we cut out at least one black point after querying $x_{i_{1}}$. For example, consider the function $g$ given in Table 1. After querying $x_{1}$, we split Table 1 into two subfunctions $g|_{x_{1}=0}$ and $g|_{x_{1}=1}$, one with $x^{(1)}$ and the other with $x^{(2)}$, since $c_{1}=(0,1)$ contains both 00 and $1$. Similarly, after querying $x_{i_{1}}$, we split the table of $f$’s black points into two non-empty tables, since $c_{i_{1}}$ is neither all-zero nor all-one. Thus, we claim $\operatorname{wt}(f_{\rho_{u,v}})\leq m-1$ for any $u,v\in\{0,1\}$.
For any $u,v\in\{0,1\}$, the number of the input variables of $f|_{\rho_{u,v}}$ is $n-3>2^{m}-3>2^{m-1}$ for any $m\geq 3$. By the induction hypothesis, we have $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{u,v}})\leq 5$. Thus, $T_{1}$’s average cost is at most
$\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}\left[\mathrm{cost}(T_{1},x)\right]\leq\frac{1}{2}\cdot 2+\Pr_{x}\left[\mathrm{cost}(T_{1},x)\geq 3\right]\cdot(3+5)=5.$
Case 2: The ECS $I$ is pure. Assume $c_{i_{1}},\ldots,c_{i_{k}}\in\{0,1\}^{m}$ are all-zero vectors without loss of generality. By Proposition 4.2, any black point $x$ must satisfy $x_{i_{1}}=\cdots=x_{i_{k}}=0$. Our query algorithm $T_{2}$ runs as follows (see Figure 1(b)):
  • (1)
    query $x_{i_{1}}$, and output 00 if $x_{i_{1}}=1$;
  • (2)
    query $x_{i_{2}}$, and output 00 if $x_{i_{2}}=1$;
  • (3)
    query $x_{j}$, and apply an optimal query algorithm to the subfunction $f|_{\rho_{0,0}}$ or $f|_{\rho_{0,1}}$.
Note that the probability of querying exactly $b\leq 2$ variables is $2^{-b}$, and that the probability of querying at least 3 variables is $1/4$.
Since $I$ is pure and maximal, the coordinate $j\not\in I$ belongs to some mixed ECS, and $c_{j}$ is neither all-zero nor all-one. Then, we have $\operatorname{wt}(f|_{\rho_{0,0}})\leq m-1$ and $\operatorname{wt}(f|_{\rho_{0,1}})\leq m-1$. Since the variable number of $\operatorname{wt}(f|_{\rho_{0,v}})$ is $n-3>2^{m}-3\geq 2^{m-1}$ for any $m\geq 3$, by the induction hypothesis, $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{0,v}})\leq 5$ for any $v\in\{0,1\}$. Thus, $T_{2}$’s average cost is at most
$\mathop{\mathbb{E}}_{x\in\{0,1\}^{n}}\left[\mathrm{cost}(T,x)\right]\leq\frac{1}{2}\cdot 1+\frac{1}{4}\cdot 2+\Pr_{x}[\mathrm{cost}(T_{2},x)\geq 3]\cdot(3+5)=3.$
Summarizing the two cases, we conclude that $\mathrm{D}_{\mathrm{ave}}(f)\leq\max\{5,3\}=5$. ∎
@@
Observe that $\mathrm{D}_{\mathrm{ave}}(f_{1}(x)\vee\cdots\vee f_{k}(x))\leq\mathrm{D}_{\mathrm{ave}}(f_{1})+\cdots+\mathrm{D}_{\mathrm{ave}}(f_{k})$ for any $f_{1},f_{2},\ldots,f_{k}$. Because one can query $f_{1}(x)$, $f_{2}(x)$, $\ldots$, $f_{k}(x)$ one by one, and compute $f_{1}(x)\vee\cdots\vee f_{k}(x)$ at the end; the expected number of variables queried is at most the sum of the expectations.
Let $k=8$. Divide the black points of $f$ evenly into $k$ disjoint parts $f_{1},\ldots,f_{k}$, each with at most $\lceil\frac{4\log n}{8}\rceil<\frac{4\log n}{8}+1$ black points. In other words, $f(x)=f_{1}(x)\vee\cdots\vee f_{k}(x)$, and $\operatorname{wt}(f_{i})\leq\lceil\frac{4\log n}{8}\rceil$.
Note that $\operatorname{wt}(f_{i})\leq\lceil\frac{4\log n}{8}\rceil\leq\frac{4\log n}{8}+1\leq\log n$ when $n\geq 4$. (When $n<4$, the lemma obviously holds.) Thus, by Lemma 4.4, $\mathrm{D}_{\mathrm{ave}}(f_{i})\leq 5$ for any $i$. Therefore, $\mathrm{D}_{\mathrm{ave}}(f)\leq\sum_{i=1}^{8}\mathrm{D}_{\mathrm{ave}}(f_{i})\leq 40$. ∎
Now, we are ready to prove Theorem 1.1.
@@
Let $m=\operatorname{wt}(f)$. If $m\leq 4\log n$, we have $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ by Corollary 4.5. Assume $m>4\log n$. Using induction on $n$, we shall prove
$\mathrm{D}_{\mathrm{ave}}(f)\leq\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+C,$ (5)
for any $n$-variable boolean function with weight $m\geq 4\cdot\log n$, where $C>0$ is a sufficiently large constant.
First, let us prove, if $m\geq n$, Lemma 4.1 immediately implies (5). Observe that $\log m+2\leq\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+C$ holds if and only if $m\geq n^{2^{-(C-2)}}\log n$. When $C\geq 4$, it is obvious that $m\geq n^{2^{-(C-2)}}\log n$. Thus, (5) follows from Lemma 4.1.
@@
  • If $\operatorname{wt}(f)\geq n$, we apply Lemma 4.1, which implies (5) as we have shown.
  • If $\operatorname{wt}(f)\leq 4\log n$, we apply the query algorithm in Corollary 4.5, which implies $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$.
  • Otherwise, query arbitrary $\ell=\lceil\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+3\rceil<\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+4$ variables, say, $x_{1},\ldots,x_{\ell}$.
  • Given the value of $x_{1},x_{2},\ldots,x_{\ell}$, say, $x_{1}=c_{1},\ldots,x_{\ell}=c_{\ell}$, apply our algorithm recursively on the subfunction $f|_{x_{1}=c_{1},\ldots,x_{\ell}=c_{\ell}}$.
Let $\rho_{1},\ldots,\rho_{2^{\ell}}$ enumerate all restrictions that fix $x_{1},\ldots,x_{\ell}$, while leaving all other variables undetermined. Let $\mathcal{R}$ be the uniform distribution over $\rho_{1},\ldots,\rho_{2^{\ell}}$. Note that $\operatorname{wt}(f)=\sum_{i=1}^{2^{\ell}}\operatorname{wt}(f|_{\rho_{i}})$, and that $\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}}[\operatorname{wt}(f|_{\rho})]=\frac{m}{2^{\ell}}\geq\frac{\log n}{16\log\frac{m}{\log n}}$. By Markov’s inequality,
$\Pr_{\rho\sim\mathcal{R}}\left[\operatorname{wt}(f|_{\rho})>4\log\frac{m}{\log n}\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}}[\operatorname{wt}(f|_{\rho})]\geq\frac{1}{4}\log n\right]\leq\frac{1}{4\log\frac{m}{\log n}}.$
According to the weight of $f|_{\rho_{i}}$, let us upper bound $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$ case by case.
Case 1: $\operatorname{wt}(f|_{\rho_{i}})\leq 4\log n$. Since $m\leq n$, we have $\ell\leq\log n+\log\log n+4$. So the number of variables in $f|_{\rho_{i}}$ is $n-\ell\geq n-(\log n+\log\log n+4)>\sqrt{n}$ (when $n$ is large enough). Notice that $\operatorname{wt}(f|_{\rho_{i}})\leq O(\log(n-\ell))$. Thus, by Corollary 4.5, we have $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})\leq C_{1}$, where $C_{1}>0$ is an absolute constant.
@@
$\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$ $\leq\log\frac{m}{\log(n-\ell)}+\log\log\frac{m}{\log(n-\ell)}+C$
$\leq 2\log\frac{m}{\log(n-\ell)}+C$
$\leq 2\log\frac{m}{\log n}+2+C,$
where the second step is because $\log\log\frac{m}{\log(n-\ell)}\leq\log\frac{m}{\log(n-\ell)}$, and the third step is because $n-\ell\geq\sqrt{n}$.
Averaging over all $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$, we have
@@
$\leq$ $\ \ell+\mathop{\mathbb{E}}_{\rho\sim\mathcal{R}}[\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})]$
$\leq$ $\ \ell+\Pr_{\rho_{i}\sim\mathcal{R}}\left[\operatorname{wt}(f|_{\rho})\leq\frac{1}{4}\log n\right]\cdot 40+\Pr_{\rho\sim\mathcal{R}}\left[\operatorname{wt}(f|_{\rho})>\frac{1}{4}\log n\right]\cdot{\left(2\log\frac{m}{\log n}+2+C\right)}$
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+4+40+\frac{1}{4\log\frac{m}{\log n}}\cdot{\left(2\log\frac{m}{\log n}+2+C\right)}$
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+45+\frac{C}{4}$
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+C,$
if $C\geq 60$.
If $m\geq 4\log n$, the right hand side of (5) is at most $\log\frac{m}{\log n}+O(\log\log\frac{m}{\log n})$, completing the proof of Theorem 1.1. ∎

### v3

@@ 3.1 Upper bound
As a warm-up, the following proposition gives the exact value of $\mathrm{D}_{\mathrm{ave}}(f)$ for boolean functions $f$ with weight $1$, such as the AND function. For convenience, we say input $x$ is a black point (with respect to $f$) if $f(x)=1$.
Proposition 3.1.
Proof.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function with a unique black point $z\in\{0,1\}^{n}$. We prove $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})$ by induction on $n$. When $n=1$, we have $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})=1$.
$=1+\frac{1}{2}\cdot 2\left(1-\frac{1}{2^{n-1}}\right)$
$=2\left(1-\frac{1}{2^{n}}\right),$
where the second step is by the induction hypothesis. ∎
Next, we show a simple bound $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+O(1)$ for any $f$. Say a query algorithm is reasonable if it terminates as soon as the subfunction becomes constant.
@@
$\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+2$ for any non-zero boolean function $f$.
@@
Let $m=\operatorname{wt}(f)$. We prove by induction on $m$ and $n$. When $m=1$, we have $\mathrm{D}_{\mathrm{ave}}(f)=2(1-\frac{1}{2^{n}})\leq 2$ by Proposition 3.1. When $n=1$, $\mathrm{D}_{\mathrm{ave}}(f)\leq 1$.
Suppose $x_{i}$ is queried first. Let $m_{0}=\operatorname{wt}(f|_{x_{i}=0})$ and $m_{1}=\operatorname{wt}(f|_{x_{i}=1})$. If $m_{b}=0$ for some $b\in\{0,1\}$, a reasonable algorithm will stop on a constant subfunction $f|_{x_{i}=b}$. Thus,
$\mathrm{D}_{\mathrm{ave}}(f)\leq 1+\frac{1}{2}(\log m+2)\leq\log m+2$
by the induction hypothesis. Otherwise, by the induction hypothesis and the AM-GM inequality, we have
@@
We introduce concepts that will be used later. Suppose $f:\{0,1\}^{n}\to\{0,1\}$ has $m$ black points $x^{(1)},\ldots,x^{(m)}\in\{0,1\}^{n}$ in lexicographical order. We call $c_{i}=(x_{i}^{(1)},\ldots,x_{i}^{(m)})$ the column pattern of coordinate $i$. Coordinates $i,j$ are positively (negatively) correlated if $c_{i}=c_{j}$ ($c_{i}=\neg c_{j}$). An equivalent coordinate set (ECS) is a set of correlated coordinates.
Say a coordinate set $S\subseteq\{1,\ldots,n\}$ is pure if each $c_{i}$ for $i\in S$ is either all-zero or all-one; otherwise, $S$ is mixed. For example, in Table 1, the set $\{5,9,{11}\}$ is pure, since $c_{5}=c_{9}=(0,0)$ and $c_{11}=(1,1)$; the set $\{1,2,3\}$ is mixed, since $c_{1}=c_{3}=(0,1)$ and $c_{2}=(1,0)$.
@@
If coordinates $i,j$ are positively (negatively) correlated, then for any $x\in\{0,1\}^{n}$ with $f(x)=1$, we have $x_{i}=x_{j}$ ($x_{i}\neq x_{j}$).
@@
Let $S$ be a mixed ECS. For any coordinate $i\in S$ and $v\in\{0,1\}$, we have $\operatorname{wt}(f|_{x_{i}=v})<\operatorname{wt}(f)$.
Proposition 3.5.
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$ be a boolean function such that $n>k\cdot 2^{\operatorname{wt}(f)-1}$, then $f$ has an ECS of size at least $k+1$.
It is straightforward to prove Propositions 3.3 and 3.4. Proposition 3.5 follows from the pigeonhole principle, since there are $2^{m-1}$ distinct equivalence classes with respect to correlation. Using these facts, one can prove that any boolean function of weight $O(\log n)$ has constant $\mathrm{D}_{\mathrm{ave}}(f)$.
@@
Let $f:\{0,1\}^{n}\rightarrow\{0,1\}$, where $\operatorname{wt}(f)<\log n$. We have $\mathrm{D}_{\mathrm{ave}}(f)\leq 5$.
@@
Let $m=\operatorname{wt}(f)\geq 3$. (Lemma 3.6 follows directly from Lemma 3.2 if $\operatorname{wt}(f)<3$.) We prove $\mathrm{D}_{\mathrm{ave}}(f)\leq 5$ by induction on $n$.
By Proposition 3.5, there exists a maximal ECS $I=\{i_{1},\ldots,i_{k}\}$ of size $k\geq 3$, since $n>2^{\operatorname{wt}(f)}=2\cdot 2^{\operatorname{wt}(f)-1}$. Without loss of generality, assume that coordinates $i_{1},\ldots,i_{k}$ are positively correlated. By Proposition 3.3, we have $x_{i_{1}}=\cdots=x_{i_{k}}$ for any black point $x\in\{0,1\}^{n}$.
If $|I|=n$, then any black point $x$ must satisfy $x_{1}=\cdots=x_{n}$. Therefore, the only possible black points are the all-zero vector and the all-one vector, so there are at most 2 black points. By Lemma 3.2, $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\operatorname{wt}(f)+2\leq 3$.
From now on, we assume $|I|<n$, and thus there exists a coordinate $j\not\in I$. Let $J$ be a maximal ECS that contains $j$. Notice that $I$ or $J$ is mixed, since at most one of $f$’s maximal ECSs can be pure.
For notational convenience, let $\rho_{u,v}$ denote the restriction fixing $x_{i_{1}}$ and $x_{i_{2}}$ to $u$, $x_{j}$ to $v$, and leaving all other variables free. Our query algorithm $T$ is defined as follows:
@@
    if $x_{i_{1}}=x_{i_{2}}$, then query $x_{j}$, and apply the query algorithm recursively on the subfunction.
The query algorithm $T$ correctly computes $f$. Input $x$ cannot be a black point if $x_{i_{1}}\neq x_{i_{2}}$, since $i_{1}$ and $i_{2}$ are positively correlated (Proposition 3.3).
For any $u,v\in\{0,1\}$, the number of inputs of $f|_{\rho_{u,v}}$ is $n-3$, and $\operatorname{wt}(f|_{\rho_{u,v}})\leq m-1$ since $I$ or $J$ is mixed (Proposition 3.4). Observe that $n-3>2^{m}-3>2^{m-1}\geq 2^{\operatorname{wt}(f|_{\rho_{u,v}})}$ for any $m\geq 3$. By the induction hypothesis, we have $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{u,v}})\leq 5$.
Let us analyze the average cost of $T$. First, notice that the probability of querying exactly 2 variables is $\frac{1}{2}$; the probability of querying at least 3 variables is $\frac{1}{2}$. Thus, we conclude that
$\mathrm{D}_{\mathrm{ave}}(f)\leq\mathop{\mathbb{E}}_{x}\left[\mathrm{cost}(T,x)\right]\leq\frac{1}{2}\cdot 2+\Pr_{x}\left[\mathrm{cost}(T,x)\geq 3\right]\cdot(3+5)=5.$
@@
We have $\mathrm{D}_{\mathrm{ave}}(f_{1}(x)\vee\cdots\vee f_{k}(x))\leq\mathrm{D}_{\mathrm{ave}}(f_{1})+\cdots+\mathrm{D}_{\mathrm{ave}}(f_{k})$ for any $f_{1},f_{2},\ldots,f_{k}$. This is because we can query $f_{1}(x)$, $f_{2}(x)$, $\ldots$, $f_{k}(x)$ one by one and compute $f_{1}(x)\vee\cdots\vee f_{k}(x)$ afterward. The expected number of variables queried is at most the sum of the individual expectations.
Let $k=8$ and $B_{f}$ denote $f$’s on-set (the set of inputs on which $f$ outputs $1$). Partition $B_{f}$ into $k$ disjoint sets $B_{1}\ldots,B_{k}$, where $|B_{i}|\leq\lceil\frac{4\log n}{8}\rceil$. Each $B_{i}$ is the on-set of some function $f_{i}$. It can be verified that $f(x)=f_{1}(x)\vee\cdots\vee f_{k}(x)$ and that $\operatorname{wt}(f_{i})=|B_{i}|\leq\lceil\frac{4\log n}{8}\rceil$. Note that $\operatorname{wt}(f_{i})\leq\lceil\frac{4\log n}{8}\rceil<\frac{1}{2}\log n+1\leq\log n$ when $n\geq 4$. (When $n<4$, the corollary holds clearly.) Thus, by Lemma 3.6, $\mathrm{D}_{\mathrm{ave}}(f_{i})\leq 5$ for any $i$, implying that $\mathrm{D}_{\mathrm{ave}}(f)\leq\sum_{i=1}^{8}\mathrm{D}_{\mathrm{ave}}(f_{i})\leq 40$. ∎
To prove Theorem 1.1, we design a query algorithm and analyze its cost. Recall that in the proof of Lemma 3.2, we considered any reasonable query algorithm, which queries an arbitrary bit and terminates until the remaining function becomes constant. Similarly, to prove Theorem 1.1, we design a more sophisticated algorithm, which queries an arbitrary variable until the subfunction satisfies the following border condition: $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ if $\operatorname{wt}(f)\leq 4\log n$ (Corollary 3.7); then the query algorithm used in the proof of Corollary 3.7 is invoked.
@@
Let $m=\operatorname{wt}(f)$. If $m\leq 4\log n$, we have $\mathrm{D}_{\mathrm{ave}}(f)=O(1)$ by Corollary 3.7. We will prove
$\mathrm{D}_{\mathrm{ave}}(f)\leq\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+87,$ (3)
by induction on $n$ for any $f:\{0,1\}^{n}\to\{0,1\}$ with $\operatorname{wt}(f)=m\geq 4\log n$.
First, when $m\geq n$, the inequality (3) directly follows from Lemma 3.2 because $\log m+2\leq\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+87$ if and only if $m\geq n^{2^{-85}}\log n$. From now on, we assume that $m\leq n$.
@@
  • (1)
    If $\operatorname{wt}(f)\geq n$, apply Lemma 3.2, which implies (3) as we have shown.
  • (2)
    Otherwise, query any
    $\ell=\left\lceil\log\frac{m}{\log n}+\log\log\frac{m}{\log n}+3\right\rceil$
    variables, say, $x_{i_{1}},\ldots,x_{i_{\ell}}$.
  • (3)
    Given the values of $x_{i_{1}},x_{i_{2}},\ldots,x_{i_{\ell}}$, say $x_{i_{1}}=c_{1},\ldots,x_{i_{\ell}}=c_{\ell}$, apply our algorithm recursively to the subfunction $f|_{x_{i_{1}}=c_{1},\ldots,x_{i_{\ell}}=c_{\ell}}$.
Let $\rho_{1},\ldots,\rho_{2^{\ell}}$ enumerate all restrictions that fix $x_{i_{1}},\ldots,x_{i_{\ell}}$, while leaving all remaining variables undetermined. Averaging over all $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$, we have
$\mathrm{D}_{\mathrm{ave}}(f)\leq\ell$ $+\mathop{\mathbb{E}}_{i}[\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})]$
$\leq\ell$ $+\Pr_{i}\left[\operatorname{wt}(f|_{\rho_{i}})\leq 4\log n\right]\cdot\mathop{\mathbb{E}}_{i}\left[\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})\mid\operatorname{wt}(f|_{\rho_{i}})\leq 4\log n\right]$
$+\Pr_{i}\left[\operatorname{wt}(f|_{\rho_{i}})>4\log n\right]\cdot\mathop{\mathbb{E}}_{i}\left[\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})\mid\operatorname{wt}(f|_{\rho_{i}})>4\log n\right].$ (4)
We have $\operatorname{wt}(f)=\sum_{i=1}^{2^{\ell}}\operatorname{wt}(f|_{\rho_{i}})$ and $\mathop{\mathbb{E}}_{i}[\operatorname{wt}(f|_{\rho_{i}})]=\frac{m}{2^{\ell}}\leq\frac{1}{8}\cdot\frac{\log n}{\log\frac{m}{\log n}}$. By Markov’s inequality,
$\Pr_{i}\left[\operatorname{wt}(f|_{\rho_{i}})>4\log n\right]\leq\frac{\mathop{\mathbb{E}}_{i}[\operatorname{wt}(f|_{\rho_{i}})]}{4\log n}\leq\frac{1}{32}\cdot\frac{1}{\log\frac{m}{\log n}}.$
We bound $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$ based on the weight of $f|_{\rho_{i}}$.
Case 1: $\operatorname{wt}(f|_{\rho_{i}})\leq 4\log n$. Since $m\leq n$, we have $\ell\leq\log n+\log\log n+4$. So the number of variables in $f|_{\rho_{i}}$ is $n-\ell\geq n-(\log n+\log\log n+4)\geq\sqrt{n}$ (when $n$ is large enough). Notice that $\operatorname{wt}(f|_{\rho_{i}})\leq 4\log n\leq 8\log(n-\ell)$. Thus, by Corollary 3.7, we have $\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})\leq 80$.
@@
$\mathrm{D}_{\mathrm{ave}}(f|_{\rho_{i}})$ $\leq\log\frac{m}{\log(n-\ell)}+\log\log\frac{m}{\log(n-\ell)}+87$
$\leq 2\log\frac{m}{\log(n-\ell)}+87$
$\leq 2\log\frac{m}{\log n}+89,$
where the second step is because $\log\log\frac{m}{\log(n-\ell)}\leq\log\frac{m}{\log(n-\ell)}$, and the third step is because $n-\ell\geq n-(\log n+\log\log n+4)\geq\sqrt{n}$.
Combining the two cases and plugging them into (4), we have
@@
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+4+80+\frac{1}{32}\cdot\frac{1}{\log\frac{m}{\log n}}\cdot{\left(2\log\frac{m}{\log n}+89\right)}$
$=$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+84+\frac{1}{16}+\frac{89}{32\cdot\log\frac{m}{\log n}}$
$\leq$ $\ \log\frac{m}{\log n}+\log\log\frac{m}{\log n}+87.$
To conclude, if $m\geq 4\log n$, the right-hand side of (3) is at most $\log\frac{m}{\log n}+O(\log\log\frac{m}{\log n})$, completing the proof of Theorem 1.1. ∎